OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (anonymous):

Use this formula: \[y-y_o=m*(x-x_o)\]and find the value of m.\[m=\frac{y-y_o}{x-x_o}\]\[m=\frac{4}{3}\]Now you can pick up the first point or the second one, I'll prefer to use the second one\[y-1=\frac{4}{3}(x-2)\]the answer is\[y=\frac{4}{3}x-\frac{5}{3}\]